# Efficacy of Image Similarity as a Metric for Augmenting Small Dataset Retinal Image Segmentation

### v1

@@ 1 Introduction
One common problem with synthetic data is how we choose to evaluate its quality (van Breugel and van der Schaar, 2023). For synthetic medical images, a common metric is the Fréchet Inception Distance (FID) (Heusel et al., 2017) which measures the similarity of two image datasets. We broadly expect more similar data to be more useful for augmentation tasks. Instinctively, we know that it is unlikely that a dataset of animal images will effectively augment a disease classification model. Similarly, we know that we can make our models more effective by adding more real data, this being very similar to the previous training data. In this paper we aim to test the validity of using the similarity, as measured by the FID, as a metric for synthetic medical images as well as deepening our understanding of the benefits and limitations of synthetic data. To do this we will specifically focus on intraretinal fluid segmentation in retinal Optical Coherence Tomography (OCT) B-scans of patients with Diabetes-related Macular Edema (DME), though the results may vary for different pathologies, image types, and segmentation tasks, as we will discuss in Section 7.
@@ 2.1 Synthetic Image Generation
A PGGAN alters this behaviour by first training the generator and discriminator at low resolutions and then, after some time, adding new, higher resolution, convolution layers. By gradually increasing the resolution over time the generator is eventually able to create higher resolution images. An example of this structure is shown in Fig. 1.
@@
In order to generate synthetic data for segmentation tasks we generate the grey-scale OCT scan in the red colour channel and the corresponding intraretinal fluid mask in the green colour channel as shown in Fig. 2 using the Zhang dataset (Kermany et al., 2018) as training data. The training dataset used has 650 OCT B-scans of the macular region of the retina across 465 patients, with 210 “NORMAL” class images, 220 images with Diabetes-related Macular Edema (DME), and 220 images with drusen. To help the model prioritise scan generation over the simpler mask generation we multiply the mask layers by a factor of 0.5. Examples of synthetic generations are shown in Appendix A.
Figure 1: Example of a PGGAN network showing the gradual addition of new high resolution layers during training (indicated as left to right). As training continues we add layers of double the size of the previous to generate images with twice the resolution. We continue this until our PGGAN is generating images with a resolution of $512\times 512$.
@@ 2.2 U-Nets
U-Nets are a common class of segmentation machine learning models that take in data of a given size and return an output with the same size as the input (Ronneberger et al., 2015). In medical imaging, the automatic generation of masks from patient data is a common application of U-Nets as they can learn to extract both superficial and more in-depth information from the images before predicting a mask over a region of interest. The name U-Net comes from the U shape of the network where images are put through two concurrent blocks of layers as shown in Fig. 3. The first half of the U-Net encodes the inputs into lower resolution and higher depth representations and the second decodes this high depth information and increases the resolution. For each deconvolution step, a second output of the corresponding convolution step is passed straight through. This allows the network to consider high resolution details after encoding and decoding steps. Learning to mask medical images is a supervised process where the prediction from the model is compared to a real mask created by an expert. Throughout training we used the unweighted binary cross entropy as a loss function to inform the network’s learning. In order to evaluate the network after training we used the Dice Similarity Coefficient (DSC) to measure the accuracy of the prediction, discussed further in the next section (Section 2.3).
@@ 2.3 Metrics
To assess the similarity of two datasets we can use the Fréchet Inception Distance (FID) (Heusel et al., 2017; Xu et al., 2018b) which is also commonly used as an equivalent for the quality of synthetic images. Given distributions of real ($\mathbb{P}_{r}$) and synthetic ($\mathbb{P}_{s}$) images, we can define an arbitrary feature function $\phi$ which gives us two output distributions $\phi(\mathbb{P}_{r})$ and ϕ(s)italic-$\phi(\mathbb{P}_{s})$ that can be compared using their means and covariances. In our case, $\phi$ is a pre-trained inception network that outputs a 2048-dimensional feature-vector for each image. This results in a 2048-dimensional distribution for both $\phi(\mathbb{P}_{r})$ and ϕ(s)italic-$\phi(\mathbb{P}_{s})$ which we can then compare using the Wasserstein-2 distance (Dowson and Landau, 1982):
@@ 3 Experimental Setup
In order to investigate the relationship between image similarity (as measured by the FID) and the segmentation U-Net model performance we propose three experiments. The first (Section 4), will validate the relationship between the number of augmentation images and performance improvements. The second (Section 5), will explore how FID impacts the relationship between the number of augmentation images and their improvements. Finally, the third experiment (Section 6), will test the relationship between the FID of the augmentation images and the improvement they provide. Each sample in these experiments is the average DSC of a U-Net model on the test dataset after it has completed training on its given training dataset. To account for random variation within model performance (and random quality of the image subset), we will take at least thirty samples per point in each experiment. The test dataset consists of 95 reserved images (28 DME, 40 Normal, 27 Drusen) that are withheld entirely from the training processes, including the PGGAN training and standard augmentation generation. In calculating the average DSC of our models we disregard images with no real or predicted mask as their DSC is undefined.
@@
Each U-Net was trained from a random initial state with the same hyper-parameters, given in Table 1 below:
@@ 4.2 Results
In Fig. 4 we can clearly see that our results broadly agree with  Frid-Adar et al. (2018) and  Anderson et al. (2022). As we add more standard images we see a significant improvement from the maximum DSC achieved before augmentation. Notably, due to the limited amount of data in our chosen augmentation method, in Fig. 4 we only see the beginnings of the plateauing seen in  Frid-Adar et al. (2018).
@@
Figure 4: The Results of Section 4 where we increase the number of images from the training data, standard augmentation data, and synthetic data. Augmentations are randomly sampled from a single synthetic dataset and a single standard augmentation technique.
@@ 5.1 Methodology
In this experiment we use the same methodology as Section 4. This time, however, we use different augmentation datasets of varied FIDs in order to explore the impacts of similarity on the shape of the relationship between the number of images and the average DSC. We expect to see higher FID datasets improve the data more slowly or not at all, whereas lower FID datasets should provide more relevant information to our training and hence have a more significant impact on the average DSC.
@@ 5.2 Results
In Fig. 5 we see that the datasets of different FIDs show varying relationships between the average DSC and the number of images in the training data. Namely we see that low FID images do not improve the model significantly, with values only slightly higher than the baseline values regardless of the number of images.
Notably in the middle plot of Fig. 5 we see that, while initially there are more significant improvements, this effect quickly reduces as we add more images at around 1,000 total images. For the rightmost plot, which has the lowest FID augmentations, we see a much more significant improvement in the average DSC as we add more images. After 1,000 total images we start to see the beginning of the expected plateauing.
@@
Figure 5: Results of Section 5 showing the mean and standard error of the mean when increasing the number of synthetic augmentations added to the training data of a U-Net model before evaluating its average DSC on the test data.
@@ 6.1 Methodology
For the blurred dataset we randomly apply Gaussian blurring to each image in our original dataset of 650 images, with a randomly sampled $\sigma$ from a uniform distribution between $\sigma_{\text{min}}$ and $\sigma_{\text{max}}$. These created datasets had FIDs shown in Table 2. Images with random noise applied had two dimensional noise drawn from a normal distribution before being multiplied by a randomly chosen scaling factor between 10 and 80. This was then added to the training data (pixel values ranging 0-255) before the image was re-normalised such that pixel values are between 0 and 255. This process yielded a dataset with an FID of 39.78. Finally, the horizontally flipped image dataset had an FID of 13.58.
@@
Similar to Section 4 we take at least 30 samples for each data-point and report the mean and the standard error of the mean.
@@ 6.2 Results
Figure 6: Results of Section 6 showing the relationship between FID and DSC after augmenting the training data with a given number of images. Here the annotation “$+N\%$” means we are adding a number of images equal to $N\%$ of the original training data size (650). Synthetic and standard augmentations follow significantly different relationships. These trends also seem to vary with the number of images added. For all cases a log-normal model is chosen to a fit a best fit line in accordance with the evidences in Table 3 but these fits are likely less reliable for lower FID values. Best fit values are given in Appendix B.
Figure 6 shows the relationship between the FID and DSC when adding varying amounts of data. We can clearly see that synthetic and standard augmentations have different relationships between their similarity and effectiveness. We also see that, broadly, synthetic data is a more effective source of augmentation data (which agrees with other studies (Alqudah et al., 2023)). However, as discussed in Section 4.2, we can improve performance using both synthetic and standard augmentations.
@@
Here the models are predicting the improvement of average DSC from the unaugmented mean DSC0. The prior probabilities for the values of these constants were chosen to try and encapsulate as much of the data-space as possible while also avoiding unrealistic results (e.g. DSC $>1$). The chosen priors are:
@@
After a U-Net model was trained with some augmentation with an FID of $x_{i}$ it was scored with an average DSC across the test dataset. This average forms one sample for that particular dataset. We then repeat this process at least 30 times for each set of augmentation images. We can then define a value $y_{i}$ as the mean of these average DSCs for each dataset which combine into $y=\{y_{1},y_{2},...,y_{n}\}$. We treat synthetic and standard augmentations as separate datasets of $x$ and $y$, as shown in the plotting of Fig. 6. We can then calculate the Bayes factor between any two given models (denoted as $M_{1}$ and $M_{2}$; in our case these will be either $f$, $g$, $h$ or $k$) by the division of their respective evidences and the model’s prior probability which in our case we treat as equal ($P(M_{1})=P(M_{2})$)(Kass and Raftery, 1995):
@@
The assumption of a normal distribution mostly matches the sampling we have seen, however each point usually has one or two outliers in their 30 samples. An example of this is shown in Fig. 7 where we took 100 samples instead of 30. In Fig. 7 we can clearly see that the Gaussian has longer tails towards low DSC values; this is likely due to the random state the U-Net started in and/or the subset of data it was augmented with. With the significantly lower sample size of 30, these outliers are often present and have a larger impact on the means of the distribution and the Bayes factor.
The log-evidences for each dataset and the models proposed in (5) are given in Table 3. These values can then be used to calculate the Bayes factor.
Synthetic Data Standard Data
@@
From Table 3 we can see that in almost every case (with the exception of Synthetic +75%) the log-normal model is preferred. In most synthetic cases it is only marginally preferred over (Or under in the case of Synthetic +75%) the horizontal line or straight line models. In Fig. 6 we therefore choose to plot a log-normal best fit line in all cases taking it as the best general model for this data. We can see that there is significant variation in the best-fit parameters (which are within the bounds of the priors) which suggests there is either no overall relationship between the improvement, the FID, and the number of images, or that there is not enough data here to adequately constrain the models. For reference, the best-fit parameters are quoted in Appendix B.
Further to this, while the model selection shows an overall preference for the log-normal model, the results suggest that future work should test synthetic datasets with significantly lower FIDs (i.e. FID $\leq 50$) to help distinguish which models are preferred, as the Bayes factors are far from definitive and may also be impacted by our choice of priors. However, this falls outside the scope of the current study, especially considering the limited data available.
The preference for log-normal distributions suggests that there may be some point along the FID-axis where images become too similar and become less effective for augmenting segmentation tasks. In some scenarios (e.g. when using different generation techniques, more training data, and/or higher baseline performance), this drop may only occur at very low FID values (FID $\leq 1$) in which case we will broadly see a significant improvement as we reduce the FID of our images. In such a scenario we could effectively use the FID as a metric for synthetic image quality for augmentation.
However, as we have seen with synthetic data in Fig. 6, if the inflection point happens at higher FID values that are more reasonably achieved with generative machine learning models (FID $\leq 10$) the use of the FID as a metric for augmentation effectiveness will be less reliable.
@@ 7 Discussion
Another consideration is that the Inception network used when calculating the FID was trained using ImageNet(Heusel et al., 2017), which does not contain medical images. This might reduce the efficacy of the FID as a measure of synthetic medical image quality. However, the results in Section 6.2 suggest there is some relationship between FID and dataset effectiveness and in Appendix A we see that the FID also follows qualitative estimates of image quality.
@@
It is also important to note that in most cases we have taken approximately 30 samples of average DSC. However, as discussed in Section 6.2 and Fig. 7, there are more outliers than one would typically expect from a normal distribution and due to the low number of samples they can have a large impact on the mean of each distribution. To help mitigate this, future studies could take more samples for each dataset but due to computational limitations this is beyond the scope of this paper.
Additionally, while the Bayes factor suggests that a log-norm relationship is more likely this preference is only substantial in one of the synthetic cases while it was significant in all standard augmentation cases. Similarly, this preference is derived from our chosen priors, and while every effort has been made to keep these reasonable and fair, it is still an important consideration.
@@ 8 Conclusion
While the results presented here are somewhat limited (see Section 7) they provide some initial insight into how we should alter our approach to metrics for synthetic data quality. Similarity, as measured by the FID, appears to be a good indicator for augmentation effectiveness especially as we add more synthetic data to our training data. However, the current evidence suggests a log-normal relationship between the FID and the DSC of our augmented model. There are a number of reasons why this might be the case. Given the small size of our dataset, the decrease in augmentation effectiveness at lower FID values is most likely due to data memorization in the PGGAN; resulting in more similar synthetic data with less novel information. These results suggest that if images are not sufficiently dissimilar they will not improve the model as much as less similar data. The FID may be lower, implying higher quality images, but in terms of the training improvement the synthetic images may not prove to be more effective than a higher FID dataset. Similarly, the results of this paper show that two different augmentation generation techniques (standard augmentations and PGGAN generated synthetic augmentations) show different relationships between similarity and augmentation effectiveness. This may also be true for different synthetic generation techniques such as diffusion models. Therefore, we suggest that new metrics are required for evaluating the effectiveness of synthetic data for augmenting machine learning tasks.

### v3

@@ 1 Introduction
One common problem with synthetic data is how we choose to evaluate its quality (van Breugel and van der Schaar, 2023). For synthetic medical images, a common metric is the Fréchet Inception Distance (FID) (Heusel et al., 2017) which measures the similarity of two image datasets. We broadly expect more similar data to be more useful for augmentation tasks. Instinctively, we know that it is unlikely that a dataset of animal images will effectively augment a disease classification model. Similarly, we know that we can make our models more effective by adding more real data, this being very similar to the previous training data. In this paper we aim to test the validity of using the similarity, as measured by the FID, as a metric for synthetic medical images as well as deepening our understanding of the benefits and limitations of synthetic data. To do this we will specifically focus on intraretinal fluid segmentation in retinal Optical Coherence Tomography (OCT) B-scans of patients with Diabetes-related Macular Edema (DME), though the results may vary for different pathologies, image types, and segmentation tasks, as we will discuss in Section˜7.
@@ 2.1 Synthetic Image Generation
A PGGAN alters this behaviour by first training the generator and discriminator at low resolutions and then, after some time, adding new, higher resolution, convolution layers. By gradually increasing the resolution over time the generator is eventually able to create higher resolution images. An example of this structure is shown in Fig.˜1.
@@
In order to generate synthetic data for segmentation tasks we generate the grey-scale OCT scan in the red colour channel and the corresponding intraretinal fluid mask in the green colour channel as shown in Fig.˜2 using the Zhang dataset (Kermany et al., 2018) as training data. The training dataset used has 650 OCT B-scans of the macular region of the retina across 465 patients, with 210 “NORMAL” class images, 220 images with Diabetes-related Macular Edema (DME), and 220 images with drusen. To help the model prioritise scan generation over the simpler mask generation we multiply the mask layers by a factor of 0.5. Examples of synthetic generations are shown in Appendix˜A.
Figure 1: Example of a PGGAN network showing the gradual addition of new high resolution layers during training (indicated as left to right). As training continues we add layers of double the size of the previous to generate images with twice the resolution. We continue this until our PGGAN is generating images with a resolution of $\times 512$ 512.
@@ 2.2 U-Nets
U-Nets are a common class of segmentation machine learning models that take in data of a given size and return an output with the same size as the input (Ronneberger et al., 2015). In medical imaging, the automatic generation of masks from patient data is a common application of U-Nets as they can learn to extract both superficial and more in-depth information from the images before predicting a mask over a region of interest. The name U-Net comes from the U shape of the network where images are put through two concurrent blocks of layers as shown in Fig.˜3. The first half of the U-Net encodes the inputs into lower resolution and higher depth representations and the second decodes this high depth information and increases the resolution. For each deconvolution step, a second output of the corresponding convolution step is passed straight through. This allows the network to consider high resolution details after encoding and decoding steps. Learning to mask medical images is a supervised process where the prediction from the model is compared to a real mask created by an expert. Throughout training we used the unweighted binary cross entropy as a loss function to inform the network’s learning. In order to evaluate the network after training we used the Dice Similarity Coefficient (DSC) to measure the accuracy of the prediction, discussed further in the next section (Section˜2.3).
@@ 2.3 Metrics
To assess the similarity of two datasets we can use the Fréchet Inception Distance (FID) (Heusel et al., 2017; Xu et al., 2018b) which is also commonly used as an equivalent for the quality of synthetic images. Given distributions of real ($\mathbb{P}_{r}$) and synthetic ($\mathbb{P}_{s}$) images, we can define an arbitrary feature function $\phi$ which gives us two output distributions $\phi(\mathbb{P}_{r})$ and $\phi(\mathbb{P}_{s})$ that can be compared using their means and covariances. In our case, $\phi$ is a pre-trained inception network that outputs a 2048-dimensional feature-vector for each image. This results in a 2048-dimensional distribution for both $\phi(\mathbb{P}_{r})$ and $\phi(\mathbb{P}_{s})$ which we can then compare using the Wasserstein-2 distance (Dowson and Landau, 1982):
@@ 3 Experimental Setup
In order to investigate the relationship between image similarity (as measured by the FID) and the segmentation U-Net model performance we propose three experiments. The first (Section˜4), will validate the relationship between the number of augmentation images and performance improvements. The second (Section˜5), will explore how FID impacts the relationship between the number of augmentation images and their improvements. Finally, the third experiment (Section˜6), will test the relationship between the FID of the augmentation images and the improvement they provide. Each sample in these experiments is the average DSC of a U-Net model on the test dataset after it has completed training on its given training dataset. To account for random variation within model performance (and random quality of the image subset), we will take at least thirty samples per point in each experiment. The test dataset consists of 95 reserved images (28 DME, 40 Normal, 27 Drusen) that are withheld entirely from the training processes, including the PGGAN training and standard augmentation generation. In calculating the average DSC of our models we disregard images with no real or predicted mask as their DSC is undefined.
@@
Each U-Net was trained from a random initial state with the same hyper-parameters, given in Table˜1 below:
@@ 4.2 Results
In Fig.˜4 we can clearly see that our results broadly agree with  Frid-Adar et al. (2018) and  Anderson et al. (2022). As we add more standard images we see a significant improvement from the maximum DSC achieved before augmentation. Notably, due to the limited amount of data in our chosen augmentation method, in Fig.˜4 we only see the beginnings of the plateauing seen in  Frid-Adar et al. (2018).
@@
Figure 4: The Results of Section˜4 where we increase the number of images from the training data, standard augmentation data, and synthetic data. Augmentations are randomly sampled from a single synthetic dataset and a single standard augmentation technique.
@@ 5.1 Methodology
In this experiment we use the same methodology as Section˜4. This time, however, we use different augmentation datasets of varied FIDs in order to explore the impacts of similarity on the shape of the relationship between the number of images and the average DSC. We expect to see higher FID datasets improve the data more slowly or not at all, whereas lower FID datasets should provide more relevant information to our training and hence have a more significant impact on the average DSC.
@@ 5.2 Results
In Fig.˜5 we see that the datasets of different FIDs show varying relationships between the average DSC and the number of images in the training data. Namely we see that low FID images do not improve the model significantly, with values only slightly higher than the baseline values regardless of the number of images.
Notably in the middle plot of Fig.˜5 we see that, while initially there are more significant improvements, this effect quickly reduces as we add more images at around 1,000 total images. For the rightmost plot, which has the lowest FID augmentations, we see a much more significant improvement in the average DSC as we add more images. After 1,000 total images we start to see the beginning of the expected plateauing.
@@
Figure 5: Results of Section˜5 showing the mean and standard error of the mean when increasing the number of synthetic augmentations added to the training data of a U-Net model before evaluating its average DSC on the test data.
@@ 6.1 Methodology
For the blurred dataset we randomly apply Gaussian blurring to each image in our original dataset of 650 images, with a randomly sampled $\sigma$ from a uniform distribution between $\sigma_{\text{min}}$ and $\sigma_{\text{max}}$. These created datasets had FIDs shown in Table˜2. Images with random noise applied had two dimensional noise drawn from a normal distribution before being multiplied by a randomly chosen scaling factor between 10 and 80. This was then added to the training data (pixel values ranging 0-255) before the image was re-normalised such that pixel values are between 0 and 255. This process yielded a dataset with an FID of 39.78. Finally, the horizontally flipped image dataset had an FID of 13.58.
@@
Similar to Section˜4 we take at least 30 samples for each data-point and report the mean and the standard error of the mean.
@@ 6.2 Results
Figure 6: Results of Section˜6 showing the relationship between FID and DSC after augmenting the training data with a given number of images. Here the annotation “$+N\%$” means we are adding a number of images equal to $N\%$ of the original training data size (650). Synthetic and standard augmentations follow significantly different relationships. These trends also seem to vary with the number of images added. For all cases a log-normal model is chosen to a fit a best fit line in accordance with the evidences in Table˜3 but these fits are likely less reliable for lower FID values. Best fit values are given in Appendix˜B.
Figure˜6 shows the relationship between the FID and DSC when adding varying amounts of data. We can clearly see that synthetic and standard augmentations have different relationships between their similarity and effectiveness. We also see that, broadly, synthetic data is a more effective source of augmentation data (which agrees with other studies (Alqudah et al., 2023)). However, as discussed in Section˜4.2, we can improve performance using both synthetic and standard augmentations.
@@
Here the models are predicting the improvement of average DSC from the unaugmented mean DSC0. The prior probabilities for the values of these constants were chosen to try and encapsulate as much of the data-space as possible while also avoiding unrealistic results (e.g. DSC >1>1> 1). The chosen priors are:
@@
After a U-Net model was trained with some augmentation with an FID of $x_{i}$ it was scored with an average DSC across the test dataset. This average forms one sample for that particular dataset. We then repeat this process at least 30 times for each set of augmentation images. We can then define a value $y_{i}$ as the mean of these average DSCs for each dataset which combine into $y=\{y_{1},y_{2},...,y_{n}\}$. We treat synthetic and standard augmentations as separate datasets of $x$ and $y$, as shown in the plotting of Fig.˜6. We can then calculate the Bayes factor between any two given models (denoted as $M_{1}$ and $M_{2}$; in our case these will be either $f$, $g$, $h$ or $k$) by the division of their respective evidences and the model’s prior probability which in our case we treat as equal ($P(M_{1})=P(M_{2})$)(Kass and Raftery, 1995):
@@
The assumption of a normal distribution mostly matches the sampling we have seen, however each point usually has one or two outliers in their 30 samples. An example of this is shown in Fig.˜7 where we took 100 samples instead of 30. In Fig.˜7 we can clearly see that the Gaussian has longer tails towards low DSC values; this is likely due to the random state the U-Net started in and/or the subset of data it was augmented with. With the significantly lower sample size of 30, these outliers are often present and have a larger impact on the means of the distribution and the Bayes factor.
The log-evidences for each dataset and the models proposed in (5) are given in Table˜3. These values can then be used to calculate the Bayes factor.
Standard Data Synthetic Data
@@
From Table˜3 we can see that in almost every case (with the exception of Standard +75%) the log-normal model is preferred. In most standard cases it is only marginally preferred over (Or under in the case of Standard +75%) the horizontal line or straight line models. In Fig.˜6 we therefore choose to plot a log-normal best fit line in all cases taking it as the best general model for this data. We can see that there is significant variation in the best-fit parameters (which are within the bounds of the priors) which suggests there is either no overall relationship between the improvement, the FID, and the number of images, or that there is not enough data here to adequately constrain the models. For reference, the best-fit parameters are quoted in Appendix˜B.
Further to this, while the model selection shows an overall preference for the log-normal model, the results suggest that future work should test synthetic datasets with significantly lower FIDs (i.e. FID 50$\leq$ 50≤ 50) to help distinguish which models are preferred, as the Bayes factors are far from definitive and may also be impacted by our choice of priors. However, this falls outside the scope of the current study, especially considering the limited data available.
The preference for log-normal distributions suggests that there may be some point along the FID-axis where images become too similar and become less effective for augmenting segmentation tasks. In some scenarios (e.g. when using different generation techniques, more training data, and/or higher baseline performance), this drop may only occur at very low FID values (FID 1$\leq$ 1≤ 1) in which case we will broadly see a significant improvement as we reduce the FID of our images. In such a scenario we could effectively use the FID as a metric for synthetic image quality for augmentation.
However, as we have seen with synthetic data in Fig.˜6, if the inflection point happens at higher FID values that are more reasonably achieved with generative machine learning models (FID 10$\leq$ 10≤ 10) the use of the FID as a metric for augmentation effectiveness will be less reliable.
@@ 7 Discussion
Another consideration is that the Inception network used when calculating the FID was trained using ImageNet(Heusel et al., 2017), which does not contain medical images. This might reduce the efficacy of the FID as a measure of synthetic medical image quality. However, the results in Section˜6.2 suggest there is some relationship between FID and dataset effectiveness and in Appendix˜A we see that the FID also follows qualitative estimates of image quality.
@@
It is also important to note that in most cases we have taken approximately 30 samples of average DSC. However, as discussed in Section˜6.2 and Fig.˜7, there are more outliers than one would typically expect from a normal distribution and due to the low number of samples they can have a large impact on the mean of each distribution. To help mitigate this, future studies could take more samples for each dataset but due to computational limitations this is beyond the scope of this paper.
Additionally, while the Bayes factor suggests that a log-norm relationship is more likely this preference is only substantial in one of the standard cases while it was significant in all synthetic augmentation cases. Similarly, this preference is derived from our chosen priors, and while every effort has been made to keep these reasonable and fair, it is still an important consideration.
@@ 8 Conclusion
While the results presented here are somewhat limited (see Section˜7) they provide some initial insight into how we should alter our approach to metrics for synthetic data quality. Similarity, as measured by the FID, appears to be a good indicator for augmentation effectiveness especially as we add more synthetic data to our training data. However, the current evidence suggests a log-normal relationship between the FID and the DSC of our augmented model. There are a number of reasons why this might be the case. Given the small size of our dataset, the decrease in augmentation effectiveness at lower FID values is most likely due to data memorization in the PGGAN; resulting in more similar synthetic data with less novel information. These results suggest that if images are not sufficiently dissimilar they will not improve the model as much as less similar data. The FID may be lower, implying higher quality images, but in terms of the training improvement the synthetic images may not prove to be more effective than a higher FID dataset. Similarly, the results of this paper show that two different augmentation generation techniques (standard augmentations and PGGAN generated synthetic augmentations) show different relationships between similarity and augmentation effectiveness. This may also be true for different synthetic generation techniques such as diffusion models. Therefore, we suggest that new metrics are required for evaluating the effectiveness of synthetic data for augmenting machine learning tasks.